# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v2

@@ 2.3. Limitations of Existing LLM-Based Solutions
A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions. Specifically, given the input token constraints of LLMs, it is often infeasible to supply an entire table and request a fully prepared table as output. Therefore, a more effective strategy employs few-shot prompting to generate data prep programs. For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”. The few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
@@
First, given the difficulties of understanding both NL questions and tables, it is challenging to determine which data prep operations are specifically required to meet the needs of each NL question, thus often resulting in false negatives (FNs) and false positives (FPs). Take the results in Figure 3 as an example: converting Age to a numerical format in code block ② is an FP, as it is irrelevant to the question. Conversely, the method fails to normalize the Date column before extracting the month in code block ③, which constitutes an FN, because it does not recognize the inconsistent Date formats.
Second, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues. For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
Recent methods, such as CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct. However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them at various granularities (table-level, column-level, and cell-level), within a single LLM agent. Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
Figure 4. A Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA methods (Right).
@@ 3. An Overview of AutoPrep
    [Uncaptioned image] Planner Agent: the Planning stage. It guides the LLM to suggest high-level data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
@@
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
@@
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and reports error messages if any code bugs occur.
After that, an [Uncaptioned image] Analyzer agent extracts the answer from the prepared table. This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper. For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4. Note that other strategies could also be used by the agent in a “plug-and-play” manner, which will be discussed in the experiments.
@@
Figure 4 illustrates how AutoPrep supports data prep for question “Which country has the most medals in total in February?” posed over a table with columns Date, Cyclist, Medal and Age.
(a) The Planning stage: The [Uncaptioned image] Planner suggests the following high-level operations to address the specific NL question:
  • $T^{\prime}$[Country]=Augment(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires aggregating the “medals in total”;
@@
  • $T^{\prime}=$Filter($T$, [Date, Country, Medal]) that filters out column Age, which is irrelevant to the question;
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, [Uncaptioned image] Augment, [Uncaptioned image] Normalizer and [Uncaptioned image] Filter, corresponding to the Augment, Normalize and Filter operations, respectively. Each specialized Programmer focuses on generating executable code for its assigned operations.
@@
After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an [Uncaptioned image] Analyzer agent to produce the answer ITA.
The Planner Agent. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
Note that our proposed AutoPrep framework is extensible. When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them. The central Planner agent can then determine which operations should be performed and assign them accordingly.
@@ 4. The Planner
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question.
@@ 4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
@@
Figure 5(a) shows an example of the aforementioned direct prompting method, which produces two high-level operations, Filter and Normalize. However, this logical plan might not be accurate as discussed as follows.
(a) Incorrect operations: The Filter operation incorrectly filters out the Cyclist column because the column is not explicitly mentioned in the question. This operation is clearly disastrous, as the country information is implicitly embedded in the Cyclist column.
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Augment operation on Cyclist is not generated, as the column has already been filtered out. Additionally, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
@@ 4.2. A Chain-of-Clauses Method
  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., Date LIKE02-%’. For each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it to prompt the LLM for generating possible high-level operations (e.g., Normalize).
@@
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f({\tt Cyclist})$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
@@
(b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows.
  • $f({\tt Cyclist})$ AS Country: this clause and relevant columns are used to prompt the LLM to generate an Augment operation.
  • ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation.
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch.
@@ 5. The Programmer & Executor
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in Figure 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
@@ 5.1. Tool-Augmented Method: Key Idea
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Augment) and an intermediate table $T_{i}$, the method generates low-level code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
Function Selection and Argument Inference. The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Aug}$) and infer the arguments (e.g., regular expression) for the selected function.
For instance, given the Augment operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Aug}$ designed specifically for Augment, obtaining an extract function with two arguments: column and func. The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation if no existing function is suitable to meet the operation’s requirements.
@@ 5.2. Search Space of Function Pools
Function pool for $\mathtt{Augment}$. We define a specific Programmer agent, i.e., Augmenter, to generate low-level code for the operation type $\mathtt{Augment}$ from the following function pool $\mathcal{F}_{\mathtt{Aug}}$.
@@
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to a lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
@@
  • $\mathtt{calculate}$. The function generates a new column through arithmetic operations of existing columns. It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns. For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [’2012’, ’2013’] and ${\tt func}$ to a lambda function lambda x: (x[’2013’]-x[’2012’])/x[’2012’].
@@
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
@@
  • $\mathtt{infer}$. This function directly uses LLMs to deduce values that could not be processed by the functions above. It takes a list of source columns and the failed target values as input, learning from demonstrations of successfully processed values to directly output the target values.
@@
Function pool for $\mathtt{Normalize}$. We define a specific Programmer agent, i.e., Normalizer, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
@@
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ as Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
@@
  • $\mathtt{format\mbox{-}datetime}$. The function standardizes the format of columns with DateTime values. It has two arguments: ${\tt column}$ and ${\tt format}$. The ${\tt format}$ argument specifies the desired format for standardization. Given the table $T_{6}$ in Figure 2(b), to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%m-%d”.
@@
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $\tt trans\_dict$ to {’Italia’: ’ITA’}.
@@
  • $\mathtt{infer}$. Similar to the one for the $\mathtt{Augment}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
@@
Function pool for $\mathtt{Filter}$. We design the Normalizer agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$. This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
Discussions. Our search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools, such as specific data prep functions for various semantic types developed in the dataprep library (Peng et al., 2021).

### v3

@@ 2.3. A Straightforward LLM-based Solution
A straightforward solution to question-aware data prep is to prompt an LLM to prepare tables, leveraging its ability to interpret the specific requirements of NL questions. For instance, consider the table in Figure 3 with columns Date, Cyclist, Medal and Age, and question: “Which country has the most medals in total in February”. A few-shot prompting strategy prompts an LLM with a task description and a few demonstrations, and requests the LLM to generate Python programs as shown in Figure 3.
@@
First, at the logical-operation level, given the inherent difficulties in understanding both NL questions and tables, it is challenging to accurately identify which data prep operations are specifically required to satisfy the needs of the NL question. This often leads to false negatives (FNs) and false positives (FPs). For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive (FP), as it is irrelevant to the question. In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative (FN), as the method ignores the inconsistency in Date formats.
Second, at the physical-operation level, due to input token limitations and challenges in long-context understanding (Li et al., 2024), it is not easy to fully understand all possible issues in a table, and thus may struggle to generate customized programs to correct issues. For example, in Figure 3, the normalization of Medal in code block ① overlooks certain corner cases (e.g., “2*”), and the country extraction in code block ⑤ fails to handle “Dav.ITA”, which is formatted differently from other values.
Recent methods, such as CoTable (Wang et al., 2024) and ReAcTable (Zhang et al., 2024), can improve few-shot prompting by employing techniques like Chain-of-Thoughts (CoT) and ReAct. However, these methods remain insufficient to tackle the challenges, as they combine all diverse tasks, such as determining operations and implementing them, within a single LLM agent. Existing studies (Ishibashi and Nishimura, 2024) have shown that a single LLM agent is often ineffective when tasked with handling a diverse range of operations, due to limited context length in LLMs and decreased inference performance with more input tokens.
Figure 4. A Multi-agent Data Prep Framework for TQA (Top), which could be plugged to any other TQA methods (Bottom).
@@ 3. An Overview of AutoPrep
    Planner Agent: the Planning stage. It guides the LLM to suggest logical data prep operations $O=\{o_{1},o_{2},\ldots,o_{|O|}\}$, which are tailored to specific question $Q$,
@@
    Multiple Programmer Agents (i.e., Normalize): the Programming stage. It directs the LLM to generate physical implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
@@
    An Executor Agent: the Executing stage. It executes the generated code and reports errors if any bugs occur.
After that, an Analyzer agent extracts the answer from the prepared table. This agent can either use LLMs as black-boxes or leverage them for code generation, which is orthogonal to the question-aware data prep problem studied in this paper. For simplicity, we use a Text-to-SQL strategy that translates the question into an SQL query over the prepared table to obtain the final answer, as shown in Figure 4. Note that other strategies could also be used by the agent in a “plug-and-play” manner.
@@
Figure 4 illustrates how AutoPrep supports data prep for an NL question posed over a table with $4$ columns.
(a) The Planning stage: The Planner suggests the following high-level operations to address the specific NL question:
  • $T^{\prime}$[Country]=Derive(“Extract country code”, Cyclist) that extracts the country information from column Cyclist, producing a new Country column, in response to the “which country” part of the question.
  • Normalize(“Case to INT”, Medal) that standardizes the value formats in the Medal column (e.g., removing quotation marks and asterisks) and then converts the strings to integers, as the question requires “the most medals”;
@@
  • $T^{\prime}$ $=$ Filter([Date, Country, Medal], $T$) that filters out column Age, which is irrelevant to the question;
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, Derive, Normalize and Filter. Each specialized Programmer focuses on generating executable code for its assigned operations.
@@
After these stages, AutoPrep generates a prepared table $T^{*}$, which is then fed into an Analyzer agent to produce the answer ITA.
The Planner Agent. The key challenge is how to suggest a logical plan that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the NL question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the method are given in Section 4.
The Programmer Agents. The key challenge is that a given logical operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Derive agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. More details of our tool-augmented approach are discussed in Section 5.
Remarks. Our proposed AutoPrep framework is extensible. When additional question-aware data prep operations are required, more specialized Programmer agents can be designed to handle them. The central Planner agent can then determine which operations should be performed and assign them accordingly.
@@ 4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Derive operation creates a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations
@@
Figure 5(a) illustrates the direct prompting method, which produces two logical operations, Filter and Normalize. However, this logical plan might not be accurate, as discussed below.
(a) Incorrect operations: The Filter operation retains the Country column simply because the question mentions “which country.” However, it fails to recognize that the original table does not actually contain a Country column. Worse yet, it incorrectly filters out the Cyclist column, merely because Cyclist is not explicitly mentioned in the question. This mistake is critical, as the country information is implicitly embedded within the Cyclist values.
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Derive operation on Cyclist is not generated, as the column has already been filtered out. Moreover, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
@@ 4.2. A Chain-of-Clauses Method
  • The second phase iteratively examines different clauses in the Analysis Sketch, e.g., ORDER BY Date. At first, we use the syntax parser from Binder (Cheng et al., 2023) to extract all UDFs and operation clauses in the sketch and sort them by the execution order. Then, for each clause, we associate the corresponding data in $T$ (e.g., values in column Date) with it and provide several in-context learning demos to prompt the LLM for generating possible logical operations (e.g., Normalize).
@@
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(A^{*}_{i},\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column $A^{*}_{i}$. Figure 5(b) shows an example Analysis Sketch with a UDF that specifies a new column Country, i.e., $f({\tt Country},{\tt Cyclist})$. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
Logical Operation Generation. Given the constructed Analysis Sketch, we orchestrate the logical operations and determine their execution order by translating the sketch into actionable steps. Specifically, we leverage the syntax parser from Binder (Cheng et al., 2023) to extract all UDFs and operation clauses from the sketch and sort them according to their intended execution sequence from the parser. Based on this parsed order, we then generate the corresponding logical operations, as illustrated below.
@@
(b) Phase II - Operation Generation: In this phase, the algorithm iteratively examines the clauses in Analysis Sketch $S$ as follows. (1) $f({\tt Country,Cyclist})$: this clause and relevant columns are used to prompt the LLM to generate an Derive operation. (2) ${\tt SUM}({\tt Medal})$: this clause and the values in column Medal are used to prompt the LLM to generate a Normalize operation. (3) Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize. Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch. Note that the Filter operation removes irrelevant columns, whereas the Derive operation creates new columns from existing ones.
@@ 5. The Programmer & Executor
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14”) or semantic formats (e.g., “ITA” and “Italia”), making it difficult to generate code customized to these tables. For example, given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021a), are needed to address specific requirements of code generation.
@@ 5.1. Tool-Augmented Method: Key Idea
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of logical operations, the method iteratively generates and executes physical executable code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$. Specifically, in the $i$-th iteration, given a high-level operation $o_{i}$ (e.g., Derive) and an intermediate table $T_{i}$, the method generates physical code and executes it to produce an updated table $T_{i+1}$ through the following two steps.
Function Selection and Argument Inference. The method first prompts the LLM to select a specific function from a function pool $\mathcal{F}$ corresponding to the operation type (e.g., $F_{\tt Der}$) and infer the arguments (e.g., regular expression) for the selected function.
For instance, given the Derive operation over table $T_{i}$ shown in Figure 6, the LLM selects a function from the pool $\mathcal{F}_{\tt Der}$ designed specifically for Derive, obtaining an extract function with two arguments: column and func. The LLM then generates preliminary code for these arguments, assigning Cyclist to the column argument and generating a lambda function with a specific regular expression for argument func.
Note that, in addition to selecting functions from the corresponding pool, our method may also prompt the LLM to write specific code for an operation (denoted as $\mathtt{CustomCode}$) if no existing function is suitable to meet the operation’s requirements.
@@ 5.2. Search Space of Function Pools
Function Pool for $\mathtt{Derive}$. We define a specific Programmer agent, i.e., Derive, to generate low-level code for the operation type $\mathtt{Derive}$ from the following function pool $\mathcal{F}_{\mathtt{Der}}$.
@@
(1) $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to a lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
@@
(2) $\mathtt{calculate}$. The function generates a new column through arithmetic operations of existing columns. It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns. For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [‘2012’, ‘2013’] and ${\tt func}$ to a lambda function lambda x: (x[‘2013’]-x[‘2012’])/x[‘2012’].
@@
(3) $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
@@
(4) $\mathtt{infer}$. This function uses LLMs to deduce values that could not be processed by the above functions. It takes a list of columns and the failed target values as input, learning from demonstrations of successfully processed values to output the target values.
@@
Function Pool for $\mathtt{Normalize}$. We define a specific Programmer agent, i.e., Normalize, to generate low-level code for the operation type $\mathtt{Normalize}$ from the following function pool $\mathcal{F}_{\mathtt{Norm}}$.
@@
(1) $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, we can use this function to standardize the Time column to integers, with ${\tt column}$ as Time and ${\tt func}$ as the lambda function.
@@
(2) $\mathtt{format\mbox{-}datetime}$. The function standardizes the format of columns with DateTime values. It has two arguments: ${\tt column}$ and ${\tt format}$. The ${\tt format}$ argument specifies the desired format for standardization. For example, to make the values in the Date column comparable, we use the $\mathtt{format\mbox{-}datetime}$ function with ${\tt column}$ set to Date and ${\tt format}$ set to “%Y-%m-%d”.
@@
(3) $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
@@
(4) $\mathtt{infer}$. Similar to the one for the $\mathtt{Derive}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
@@
Function Pool for Filter. We design the Filter agent to utilize a function pool $\mathcal{F}_{\mathtt{Filter}}$ with a single function, $\mathtt{filter\mbox{-}columns}$. This function has one argument, ${\tt rel\_columns}$, which represents a list of question-related column names.
Remarks. The design of $\mathtt{CustomCode}$ and $\mathtt{infer}$ operations enables AutoPrep to generalize to unseen tasks. For example, when encountering missing or clearly erroneous values, AutoPrep can directly prompt an LLM to infer plausible replacements. Moreover, our search space is extensible: additional functions or external APIs can be easily integrated into the function pool, including specialized data preparation functions for various semantic types, such as those provided by the dataprep library (Peng et al., 2021a).